OpenStudy (rikkie):

32^3/5 in simplest form?

OpenStudy (anonymous):

What do you mean? \[32^{3/5}\] or \[\frac{32^{3}}{ 5 }\]

OpenStudy (rikkie):

The first!

OpenStudy (anonymous):

okay, this will end up as a radical.

OpenStudy (rikkie):

5√^3?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

(Sorry about delayed response. For some reason the page would not load.)

OpenStudy (anonymous):

\[\sqrt[5]{x^3}\]

OpenStudy (anonymous):

or in this case 32=x

OpenStudy (rikkie):

That's alright, it was the same for me. I'm stuck on the next step and I have it like 5√2*5 5v2*5 5v2*5 but don't know if that's the right track

OpenStudy (mathmale):

Note: 32 is a perfect 5th power, so it should be easier to evaluate 32^(3/5) if you re-write the whole problem as\[(2^5)^\frac{ 3 }{ 5 }\]

OpenStudy (mathmale):

What do rules of exponents tell you about evaluating \[(a^b)^c\]?

OpenStudy (mathmale):

\[(a^b)^c=?\]

OpenStudy (rikkie):

Do you add the b and c?

OpenStudy (mathmale):

\[(2^5)^{3/5}=?\]

OpenStudy (mathmale):

No, you don't add b and c. Strongly suggest that you look up "rules of exponentiation" and copy them down for later reference.

OpenStudy (rikkie):

I'm still fuzzy with these, and I'll remember that!

OpenStudy (mathmale):

\[(a^c)(a^d)=a ^{c+d},\] but this is a different situation from the one you're deaing with.

OpenStudy (mathmale):

\[(a^b)^c=?\]

OpenStudy (rikkie):

I'm still quite confused, could you explain a bit?